OpenStudy (anonymous):

Trevor, currently 22 years old, is half of his dad's age six years ago. How old is Trevor's Dad? Answer 40 44 50 54 Trevor, currently 22 years old, is half of his dad's age six years ago. How old is Trevor's Dad? Answer 40 44 50 54 @Mathematics

OpenStudy (anonymous):

twice of what ???

OpenStudy (diyadiya):

Trevor's age = 22 Trevor's Dad's age = 'd' (22) = 1/2(d - 6) 44 = d - 6 dad's age = 50

OpenStudy (anonymous):

see things are simple u just need to unlock them see how diya unlocked it right nothing is tough it is we who makes things look tough

OpenStudy (anonymous):

so buddy b4 puttin down ur weapons at least give a try to it!@!!!